OpenStudy (anonymous):

what's the difference between those methods of finding the average , for 8 numbers 98,77,86,92,80,70,94,65 method 1 :- sumation/8 method 2 :- 98+77/2=87.5 87.5+86/2=86.75 .... and so on avg1=82.75 avg2=75.3359..

OpenStudy (kirbykirby):

Method one is correct. But if you do method 2: \[ \frac{98+77}{2}=87.5\\ \, \\ \frac{87.5+86}{2}=86.75\\ ...\] Essentially what you are doing is this: \[\Large \frac{\frac{\frac{98+77}{2}+86}{2}+92}{2}...\]

OpenStudy (anonymous):

then it depand on the arrange of numbers and get diffrent answer each time right ?

OpenStudy (kirbykirby):

Yeah

OpenStudy (anonymous):

so what if i arrange them from the smallest ?

OpenStudy (anonymous):

wont that be good approximation for the avg ?

OpenStudy (kirbykirby):

You mean if you do: \[\Large \frac{\frac{\frac{65+70}{2}+77}{2}+80}{2}...\] instead?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i would get 94.13 which is far away from avg . hmm so which is better arrange them randomly ?

OpenStudy (kirbykirby):

Well method 2 is simply an incorrect method of finding the mean.

OpenStudy (anonymous):

naaa lol dnt say its wrong haha im doing my senior project to convince ppl it culd be approximation method :P its not wrong 100%

OpenStudy (kirbykirby):

Well I suppose there are other ways to calculate the mean, like the geometric mean or the harmonic mean. I'm not familiar with this method of calculating it. And the fact that it gives different answers when you order the numbers differently doesn't seem terribly useful

OpenStudy (kirbykirby):

But if you can do a proof for it, that's better than just giving examples.

OpenStudy (anonymous):

ik :D im just doing my jop its worth if it give good approximation for n > 2000 or somthing

OpenStudy (kirbykirby):

maybe. DId you prove that?

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

i think if the numbers were ordered and the difference between each 2 adjacent numbers remained the same then you could use method 2, right?

OpenStudy (anonymous):

that would be complicated , but some how uve got a ponit i have to determine the Error range